OpenStudy (anonymous):

The height in feet above the ground of a ball thrown upwards from the top of a building is given by s = -16x^2 + 160x + 200, where t is the time in seconds a) what is the maximum height (in feet) of the ball? b) what is v^-(32)

OpenStudy (agentnao):

Use a graphing calculator to solve this. It'll be the maximum point of the graph.

OpenStudy (anonymous):

use vertex formula -b/2a

OpenStudy (anonymous):

160/2(32) 160/64

OpenStudy (agentnao):

Should be 600

OpenStudy (anonymous):

use vertex for which problem A or B

OpenStudy (phi):

Your equation uses x, but I assume it should be t. so t= -b/2a, and you must plug this value into the equation to find the maximum height.

OpenStudy (anonymous):

what about B

OpenStudy (phi):

what is v^(-32)?

OpenStudy (anonymous):

yes

OpenStudy (agentnao):

-b/2a doesn't really work for this number. The height of the top of a building won't be -160/32...

OpenStudy (anonymous):

-b/2a 160/64=2.5 plug in 2.5 into your equation

OpenStudy (anonymous):

so you take derivative of equation first?

OpenStudy (phi):

is v velocity? then yes, take the derivative.

OpenStudy (anonymous):

no i mean for maximum height of ball

OpenStudy (phi):

t= -b/2a = 5 s = -16*5^2 + 160*5 + 200= 600

OpenStudy (phi):

Or take the derivate, set it = 0, and solve for t. You get t=5

OpenStudy (anonymous):

What about B?????